# Natural polynomials for Kerr quasi-normal modes

### v2

@@ III Preliminaries
–  In many areas of physics, polynomial solutions to a system’s equations of motion play an important role in describing that system’s natural modes Sakurai (1993); Cook and Zalutskiy (2014); London (2023b); Fackerell and Crossman (2008). For the QNM’s radial problem (Eq. III), one approach to understanding polynomial solutions begins by studying the differential part of $\mathcal{L}_{\xi}$, namely
@@ IV The radial problem with Confluent Heun Polynomials
When $\alpha_{p}\neq 0$, the off-diagonal elements of $\hat{L}{{}_{p}}^{(\text{cH})}$ will be proportional to ${\langle{y_{pk^{\prime}}}\,|\,{\xi\,|\,y_{pk}}\rangle}$. It was found in Paper I that $|{\langle{y_{pk^{\prime}}}\,|\,{\xi\,|\,y_{pk}}\rangle}|$ do not typically decrease in value as $|k^{\prime}-k|$ increases. Consequently, $\hat{L}{{}_{p}}^{(\text{cH})}$ will typically not be non-band-diagonal when $\alpha_{p}\neq 0$.
@@
–  From inspection of Eq. (30a), it is clear that when $\alpha_{p}=0$, $\hat{L}{{}_{p}}^{(\text{cH})}$ is diagonal. This implies that $\hat{L}{{}_{p}}^{(\text{cH})}$ may be approximately diagonal when $|\alpha_{p}|\ll 1$.
@@
In Eq. (33), since ${p_{\star}^{+}}$ and ${p_{\star}^{-}}$ always come as a pair, they are duals of each other. Henceforth, ${p_{\star}^{+}}$ wll be referred to as the upper pseudo-polynomial order, and ${p_{\star}^{-}}$ will be referred to as the lower pseudo-polynomial order.
$n$
$\quad\quad\quad\quad\quad{p_{\star}^{+}}$
$\quad\quad\quad\quad\quad{p_{\star}^{-}}$
00
$\quad 1.3879+0.5964i$
$\quad-0.6768+2.1304i$
$1$
$\quad 2.1725+0.5415i$
$\quad-0.0230+2.0846i$
$2$
$\quad 2.9791+0.4394i$
$\quad 0.6490+1.9996i$
$4$
$\quad 4.6187+0.1537i$
$\quad 2.0152+1.7615i$
$6$
$\quad 6.2586+0.0743i$
$\quad 3.3816+1.6954i$
$8$
$\quad 15.2263+-1.9170i$
$\quad 10.8539+0.0362i$
Table 1: Select pseudo-polynomial orders, ${p_{\star}^{+}}$ and ${p_{\star}^{-}}$, for BH spin of $a/M=0.7$, and QNM indices $(s,\ell,m)=(-2,2,2)$. Parameters for select QNM overtone numbers, $n$, are shown.
For the moment, we are concerned with the implications of this duality on solutions to the radial problem (Eq. IV), the QNMs radial functions, and the ultimate practicality of confluent Heun polynomial representations for non-polynomial problems.
@@
It also follows that ${\text{C}_{1}}=-{p_{\star}^{+}}{p_{\star}^{-}}$. In terms of physical parameters (Eq. III), ${p_{\star}^{+}}$ and ${p_{\star}^{-}}$ are
[…]
@@
From Eq. (34), it may be seen that if ${p_{\star}^{+}}$ is a polynomial order, then ${p_{\star}^{-}}$ typically will not be unless ${\text{C}_{4}}$ is a sufficiently negative integer. In such cases, the solution space of Eq. (IV) will be comprised of ${p_{\star}^{+}}+1$ polynomial solutions, and an infinity of non-polynomial solutions. In this way, there exists a confluent Heun polynomial/non-polynomial duality.
For QNMs, it is expected that neither ${p_{\star}^{+}}$ nor ${p_{\star}^{-}}$ will be polynomial powers, and that the proximity of each to an integer will vary with different values of the overtone index, $n$. In particular, solutions to the QNM radial problem may only be closely associated with confluent Heun polynomials when $\tilde{\omega}_{{\ell mn}}$ are such that ${p_{\star}^{+}}$ or ${p_{\star}^{-}}$ is close to a positive integer in the sense that,
$|\text{Im}\,{p_{\star}^{\pm}}|\;\ll\;\text{Re}\,{p_{\star}^{\pm}}\;,$ (37a)
@@
${p_{\star}^{\pm}}-\lfloor\text{Re}\,{p_{\star}^{\pm}}\rfloor\;\ll\;1\;.$ (37b)
In Eq. (IV), $\lfloor x\rfloor$ is the integral part of a real number (i.e. $\lfloor 3.1416\rfloor=3$).
Table (1) provides examples of ${p_{\star}^{\pm}}$ for select QNM cases with mode numbers $(\ell,m)=(2,2)$. There it is observed that the QNM’s ${p_{\star}^{+}}$ only very roughly tracks the overtone index, being approximately off by one for e.g. $n=2$, but off by significantly less than one for $n=6$. Notably, $n=8$ departs from this trend, having a $\text{Re}\,{p_{\star}^{+}}$ that is just under two times the overtone index. In all cases, it is observed that $n<{p_{\star}^{+}}$. Further, the case of the fundamental QNM, $n=0$, demonstrates a general observation, namely that the $n=0$ QNMs are typically not close to confluent Heun polynomial systems in that their ${p_{\star}^{\pm}}$ are not close to integers.
@@
–  Since the QNMs ${p_{\star}^{\pm}}$ are typically complex valued, it is not guaranteed that using an order-$p$ confluent Heun polynomial basis with e.g. $p=\lfloor\mathrm{max}(|{p_{\star}^{+}}|,|{p_{\star}^{-}}|)\rfloor$ will yield an accurate representation of the radial problem. Worse, simply increasing the order of the confluent Heun polynomial basis (i.e. increasing $p$) does not necessarily result in a more precise representation. This may be understood as resulting from a basic limitation of approximation with a fixed-order polynomial bases.
@@
[…]
Clearly, in this case it would be wise to represent $|{1}\rangle$ with the order-00 confluent Heun polynomial. This is consistent with the completeness of each confluent Heun polynomial fixed order basis: an order-$p$ confluent Heun polynomial basis can exactly represent polynomials of the same order; while lower order polynomials may also be represented, their representation in polynomial bases of much larger order increases the likelihood of numerical noise.
@@
In contrast, a basis of $P+1$ classical polynomials (i.e. with unique orders ranging from 00 to $P$) would only have $\frac{1}{2}(P+1)^{2}$ coefficients to total. More importantly, unlike the order-$P$ confluent Heun polynomials, such a basis would have expansion coefficients (analogs of ${\langle{y_{Pk}}\,|\,{f}\rangle}$) that necessarily decrease as $k$ increases, meaning that fewer than $\frac{1}{2}(P+1)^{2}$ coefficients may ultimately be of practical importance.
@@ V Natural & Canonical Polynomials
[…]
@@
$c_{20}{\langle{u_{0}}\,|\,{y_{20}}\rangle}+c_{21}{\langle{u_{0}}\,|\,{y_{21}}\rangle}+{\langle{u_{0}}\,|\,{y_{22}}\rangle}\;$ = 0,absent 0=\;0\;,= 0 , (45a)
$c_{20}{\langle{u_{1}}\,|\,{y_{20}}\rangle}+c_{21}{\langle{u_{1}}\,|\,{y_{21}}\rangle}+{\langle{u_{1}}\,|\,{y_{22}}\rangle}\;$ = 0.absent 0=\;0\;.= 0 . (45b)
@@
Equation (V) is equivalent to the linear matrix equation,
@@
In Eqs. (V) and (47), the essential features of subsequent cases (i.e. $p>2$) are present: For any $|{u_{p}}\rangle$, there will be $p$ equations of the form of Eq. (V), with a total of $p$ unknowns, and this system of linear equations will generally be equivalent to a linear matrix equation,
@@
Since the scalar product (Eq. 14) is unchanged with respect to negation of both of its inputs (e.g. ${\langle{u_{p}}\,|\,{u_{p}}\rangle}={\langle{-u_{p}}\,|\,{-u_{p}}\rangle}$), there is an overall ±$\pm 1$ scaling, or rather “signature” freedom. Here, this is fixed by dividing each $u_{p}$ by its value at $\xi=1$,
@@
Similarly, the canonical confluent Heun polynomials may be defined by applying Gram-Schmidt to the monomials using the weight function, Eq. (16). Orthogonality is guaranteed because the algorithm iteratively removes the non-orthogonal components between different elements: In particular, given a sequence of monomials $\{\xi^{p}\}_{p=0}^{\infty}$, the algorithm iteratively defines canonical polynomials, $u_{p}(\xi)$, to be $\xi^{p}$ plus a lower order polynomial that is constrained by requiring that ${\langle{u_{p}}\,|\,{u_{p^{\prime}}}\rangle}=0$ for all $p^{\prime}<p$. The results is a sequence of orthogonal polynomials that can then be normalized according to Eq. (V).
It happens that that the weight function used to define the scalar product in Eq. (14) is of the Pollaczek-Jacobi type, but with complex valued parameters (Eqs. III and III). Since our weight function may generally be evaluated using analytic continuation (see Eq. III), the Gram-Schmidt polynomials described above may be thought of as analytically continued Pollaczek-Jacobi polynomials Chen and Dai (2010). Further connections between the polynomials described above, and the Pollaczek-Jacobi polynomials will be explored in future work Gurevich and London (2024).
@@
$u_{1}(\xi)\;$ $=\;\left[{\frac{{\langle\xi^{0}\rangle}}{{\langle\xi^{0}\rangle\langle\xi^{2}\rangle-\langle\xi^{1}\rangle^{2}}}}\right]^{1/2}\;\left(\frac{\langle\xi^{1}\rangle}{\langle\xi^{0}\rangle}+\xi\right)\;.$ (51b)
In Eq. (V), we recall that $\langle\xi^{k}\rangle$ are monomial moments defined by Eq. (III).
Figure 2: The first five canonical confluent Heun polynomials, i.e. $u_{p}(\xi)$ with $p\in\{0,1,2,3,4\}$. Polynomials are constructed using field spin weight $s=-2$, BH spin $a/M=0.7$, and Kerr QNM frequency $M\tilde{\omega}_{220}=0.5326-0.0808i$. For ease of presentation, polynomials have been divided by their value at $\xi=1$ (i.e. spatial infinity). Top panel: Real part of each polynomial. Bottom panel: imaginary part of each polynomial. Related discussion is provided in Sec. VII.
On arrives at Eqs. (51a) and (51b) regardless of whether direct or Gram-Schmidt construction is used.
@@
In practice, since the confluent Heun polynomials are eigenfunctions to $\mathcal{D}{{}_{\xi}}$, it is useful to represent each $|{u_{p}}\rangle$ in terms of confluent Heun polynomials, $|{y_{p^{\prime}k^{\prime}}}\rangle$. However, since the confluent Heun polynomials are overcomplete, there is no unique representation.
@@
Within Eq. (V), each canonical confluent Heun polynomial is represented in the basis of a different polynomial order, meaning that a set of multiple canonical confluent Heun polynomials would require the generation of confluent Heun polynomials from multiple orders. For this reason, Eq. (V) will henceforth be referred to as the multi-order representation of the canonical confluent Heun polynomials.
@@
In Eqs. (54a) and (54b), the identity operator for the order-$p^{\prime}$ confluent Heun polynomials has been applied to $|{u_{p}}\rangle$, and in Eq. (54c), $d_{p^{\prime}pk}={\langle{y_{p^{\prime}k}}\,|\,{u_{p}}\rangle}$.
$d_{p^{\prime}pk}\;=\;{\langle{y_{p^{\prime}k}}\,|\,{u_{p}}\rangle}$ (55)
In practice, one may first generate $|{u_{p}}\rangle$ via Gram-Schmidt construction, and then explicitly compute ${\langle{y_{p^{\prime}k}}\,|\,{u_{p}}\rangle}$ using Eq. (19). Equation (V) will henceforth be referred to as the single-order representation of the canonical confluent Heun polynomials. If the largest polynomial order of interest is $p^{\prime}$, then the single-order representation has the practical advantage of only requiring $p^{\prime}+1$ confluent Heun polynomials to be computed after Gram-Schmidt construction.

### v3

@@ III Preliminaries
–  In many areas of physics, polynomial solutions to a system’s equations of motion play an important role in describing that system’s natural modes Sakurai (1993); Cook and Zalutskiy (2014); London (2023b); Fackerell and Crossman (2008). For the QNM’s radial problem (Eq. 13), one approach to understanding polynomial solutions begins by studying the differential part of $\mathcal{L}_{\xi}$, namely
@@
Lastly, the overcompleteness of the confluent Heun polynomails does not imply that it is trivial to use them as a basis for general solutions to the QNM radial problem. Doing so presents two challenges that we will discuss in upcoming sections. The first challenge is that each fixed-order confluent Heun polynomial basis need not have well ordered expansion coefficients, meaning that there is not necessarily an order to how much each polynomial contributes to a radial expansion. This challenge will be discussed in the upcoming section, Sec. IV.
The second challenge is that the matrix representation of e.g. $\mathcal{L}_{\xi}$ in a finite dimensional basis need not have exactly the same spectra as its infinite dimensional counterpart (i.e. there may appear modes that are purely the effect of truncation) Petropoulou and Velázquez (2014). This challenge will be revisited in Sec. VII.
@@ IV The radial problem with Confluent Heun Polynomials
When $\alpha_{p}\neq 0$, the off-diagonal elements of $\hat{L}{{}_{p}}^{(\text{cH})}$ will be proportional to $\langle{y_{pk^{\prime}}}\,|\,{\xi\,|\,y_{pk}}\rangle$}. It was found in Paper I that $|{\langle{y_{pk^{\prime}}}\,|\,{\xi\,|\,y_{pk}}\rangle}|$ do not typically decrease in value as $|k^{\prime}-k|$ increases. Consequently, $\hat{L}{{}_{p}}^{(\text{cH})}$ will typically be neither band-diagonal nor approximately diagonal when $\alpha_{p}\neq 0$.
@@
–  From inspection of Eq. (30a), it is clear that when $\alpha_{p}=0$, $\hat{L}{{}_{p}}^{(\text{cH})}$ is diagonal. This implies that $\hat{L}{{}_{p}}^{(\text{cH})}$ may be approximately diagonal when $|\alpha_{p}|\ll 1$. In the case where $\alpha_{p}=0$, a finite subset of the solution space will be comprosed of polynomails, and the remainder will be an infinite set of non-polynomial functions.
It happens that, when $\alpha_{p}\neq 0$, the solution space will typically be comprised of two qualitatively destict sectors: some solutions will be approximately polynomial in the sense that their series expansions converge quickly after some particular term, and other strongly non-polynomialsolutions whose series expansions converge relatively slowly (see e.g. the asymptotic analysis in Ref. Nollert (1999)). Henceforth, we refer to the presence of these two solutions types as “polynomial / non-polynomial duality [^6] We expect “polynomial / non-polynomial duality” to be related to the existence of quasi-bound states for $s=0$ Brito et al. (2015); Vieira et al. (2022).”.
@@
In what follows, $p_{\star}^{+}$} wll be referred to as the upper pseudo-polynomial order, and $p_{\star}^{-}$} will be referred to as the lower pseudo-polynomial order.
$n$ $p_{\star}^{+}$} $p_{\star}^{-}$}
0 $1.3879+0.5964i$ 0.6768+2.1304i\!\!\!\!-0.6768+2.1304i
11 $2.1725+0.5415i$ 0.0230+2.0846i\!\!\!\!-0.0230+2.0846i
22 $2.9791+0.4394i$ $0.6490+1.9996i$
44 $4.6187+0.1537i$ $2.0152+1.7615i$
66 $6.2586+0.0743i$ $3.3816+1.6954i$
88
$10$ $11.0503+0.0551i$ $7.3742+1.6794i$
$12$ $13.0252+0.0344i$ $9.0198+1.6621i$
$14$ $15.0134+0.0152i$ $10.6765+1.6461i$
$16$ 17.01230.0004i17.0123-0.0004i $12.3420+1.6331i$
$18$ 19.02040.0084i19.0204-0.0084i $14.0152+1.6265i$
2020 21.02980.0007i21.0298-0.0007i $15.6895+1.6329i$
Table 1: Select pseudo-polynomial orders, $p_{\star}^{+}$} and $p_{\star}^{-}$}, for BH spin of $a/M=0.7$, and QNM indices $(s,\ell,m)=(-2,2,2)$. Parameters for select QNM overtone numbers, $n$, are shown. Only even $n$ are shown for brevity, and results for odd $n$ show the same qualitative trend (e.g. for $n=13$, $({p_{\star}^{+}},{p_{\star}^{-}})=(14.0178+0.0245i,9.8469+1.6539i)$). We do not show the result for $n=8$, where further conditions are needed due to the appearance of non-traditional modes Cook and Zalutskiy (2014).
For the moment, we are concerned with the implications of the pseudo-polynomial orders on solutions to the radial problem (Eq. 28), the QNMs radial functions, and the ultimate practicality of confluent Heun polynomial representations for non-polynomial problems.
@@
It also follows that ${\text{C}_{1}}=-{p_{\star}^{+}}{p_{\star}^{-}}$. In terms of physical parameters (Eq. 12), $p_{\star}^{+}$} and $p_{\star}^{-}$} are
${p_{\star}^{+}}\;$ $=\;-1\,+\,4iM\tilde{\omega}\;$, (35a)
${p_{\star}^{-}}\;$ $=\;{p_{\star}^{+}}\,-s\,-\,2iM\tilde{\omega}\,-\,\frac{i\left(am-2M^{2}\tilde{\omega}\right)}{\delta}\;.$ (35b)
In Eq. (35) we have simplified Eq. (33) by adopting the strategy of Ref. Teukolsky (1973), and holding that if $x$ is a complex number, then e.g. $\sqrt{-x^{2}}=i\,x$, rather than the generally correct value of $i\,x\,\text{sign}{(\text{Im}(x))}$. The effect of this choice is to swap $p_{\star}^{+}$} with $p_{\star}^{-}$}. Thus, it should be understood that Eqs. (35) and (33) are formally different.
@@
From Eq. (34), three scenarios may be deduced for which the solution space includes a finite number of polynomials in addition to an infinity of non-polynomials:
  • (i)
    The first scenario is that, if $p_{\star}^{+}$} is a positive integer, then $p_{\star}^{-}$} typically will not be unless $\text{C}_{4}$} is a sufficiently negative integer.
  • (ii)
    The second scenario is simply that of the first, but with $p_{\star}^{-}$} being a positive integer, and $p_{\star}^{-}$} having a fractional part.
  • (iii)
    In the remaining scenario both $p_{\star}^{-}$} and $p_{\star}^{+}$} are integers. In this case, it is possible for the solution space to support two sets of polynomials, each of different order, as well as an infinite number of non-polynomial solutions.
Generally, neither $p_{\star}^{+}$} nor $p_{\star}^{-}$} will be positive integers for the QNMs, and their proximity to a positive integer will vary with different values of the overtone index, $n$. In particular, solutions to the QNM radial problem may only be closely associated with confluent Heun polynomials when $\tilde{\omega}_{{\ell mn}}$ are such that $p_{\star}^{+}$} or $p_{\star}^{-}$} is close to a positive integer according to
${|\text{Im}\,{p_{\star}^{\pm}}}/{\text{Re}\,{p_{\star}^{\pm}}}|\;\ll\;1\;,$ (37a)
@@
${p_{\star}^{\pm}}-\left[\text{Re}\,{p_{\star}^{\pm}}\right]\;\ll\;1\;.$ (37b)
In Eq. (37), $\left[x\right]$ denotes rounding to the nearest integer (e.g. $\left[3.1416\right]=3$).
Table (1) provides examples of $p_{\star}^{\pm}$} for select QNM cases with mode numbers $(\ell,m)=(2,2)$. Values of $p_{\star}^{\pm}$} have been defined using Eq. (33). It is observed that the QNM’s $p_{\star}^{+}$} only very roughly tracks the overtone index, being approximately off by one for e.g. $n=2$, but off by significantly less than one for $n=6$ and above. The case of the fundamental QNM, $n=0$, demonstrates a general observation, namely that the $n=0$ QNMs are typically not close to confluent Heun polynomial systems in that their $p_{\star}^{\pm}$} are not close to integers.
@@
–  Since the QNMs $p_{\star}^{\pm}$} are typically complex valued, it is not guaranteed that using an order-$p$ confluent Heun polynomial basis with e.g. $p=\lfloor\mathrm{max}(|{p_{\star}^{+}}|,|{p_{\star}^{-}}|)\rfloor$ will yield an accurate representation of the radial problem. Worse, simply increasing the order of the confluent Heun polynomial basis (i.e. increasing $p$) does not necessarily result in a more precise representation. This may be understood as resulting from a basic limitation of approximation with fixed-order polynomial bases.
@@
$|{1}\rangle\;=\;\sum_{k=0}^{50}\,|{y_{50,k}}\rangle{\langle{y_{50,k}}\,|\,{1}\rangle}\;.$ (38)
Each polynomial, $|{y_{50,k}}\rangle$, has $51$ coefficients, and on the right hand side of Eq. (38), values of $\langle{y_{50,k}}\,|\,{1}\rangle$} must conspire such that only constant terms survive the sum (as is required since $1$ is a constant). In practice, finite precision means that the cancellations will not be exact. As a result, the net approximation will be contaminated by high frequency polynomial oscillations, or rather, polynomial noise.
Clearly, in this case it would be wise to represent $|{1}\rangle$ with the order-0 confluent Heun polynomial. This is consistent with the completeness of each confluent Heun polynomial fixed order basis: an order-$p$ confluent Heun polynomial basis can exactly represent polynomials of the same order; while lower order polynomials may also be represented, their representation in polynomial bases of much larger order increases the likelihood of numerical noise.
@@
In contrast, a basis of $P+1$ classical polynomials (i.e. with unique orders ranging from 0 to $P$) would only have $\frac{1}{2}(P+1)^{2}$ coefficients to total. More importantly, unlike the order-$P$ confluent Heun polynomials, such a basis would have expansion coefficients (analogs of $\langle{y_{Pk}}\,|\,{f}\rangle$}) that necessarily decrease as $k$ increases, meaning that fewer than $\frac{1}{2}(P+1)^{2}$ coefficients may ultimately be of practical importance.
@@ V Natural & Canonical Polynomials
${\langle{u_{0}}\,|\,{u_{1}}\rangle}\;=\;c_{10}{\langle{y_{00}}\,|\,{y_{10}}\rangle}+{\langle{y_{00}}\,|\,{y_{11}}\rangle}\;=\;0\;.$ (43)
@@
$c_{20}{\langle{u_{0}}\,|\,{y_{20}}\rangle}+c_{21}{\langle{u_{0}}\,|\,{y_{21}}\rangle}+{\langle{u_{0}}\,|\,{y_{22}}\rangle}\;$ = $=\;0\;,$ (45a)
$c_{20}{\langle{u_{1}}\,|\,{y_{20}}\rangle}+c_{21}{\langle{u_{1}}\,|\,{y_{21}}\rangle}+{\langle{u_{1}}\,|\,{y_{22}}\rangle}\;$ = $=\;0\;.$ (45b)
@@
Equation (46) is equivalent to the linear matrix equation,
@@
In Eqs. (46) and (47), the essential features of subsequent cases (i.e. $p>2$) are present: For any $|{u_{p}}\rangle$, there will be $p$ equations of the form of Eq. (46), with a total of $p$ unknowns, and this system of linear equations will generally be equivalent to a linear matrix equation,
@@
Since the scalar product (Eq. 14) is unchanged with respect to negation of both of its inputs (e.g. ${\langle{u_{p}}\,|\,{u_{p}}\rangle}={\langle{-u_{p}}\,|\,{-u_{p}}\rangle}$), there is an overall $\pm 1$ scaling, or rather “signature” freedom. Here, this is fixed by dividing each $u_{p}$ by its value at $\xi=1$,
@@
Similarly, the canonical confluent Heun polynomials may be defined by applying Gram-Schmidt to the monomials using the weight function, Eq. (16). Orthogonality is guaranteed because the algorithm iteratively removes the non-orthogonal components between different elements:
Figure 2: The first five canonical confluent Heun polynomials, i.e. $u_{p}(\xi)$ with $p\in\{0,1,2,3,4\}$. Polynomials are constructed using field spin weight s=2s=-2, BH spin $a/M=0.7$, and Kerr QNM frequency $M\tilde{\omega}_{220}=0.5326-0.0808i$. For ease of presentation, polynomials have been divided by their value at $\xi=1$ (i.e. spatial infinity). Top panel: Real part of each polynomial. Bottom panel: imaginary part of each polynomial. Related discussion is provided in Sec. VII.
In particular, given a sequence of monomials $\{\xi^{p}\}_{p=0}^{\infty}$, the algorithm iteratively defines canonical polynomials, $u_{p}(\xi)$, to be $\xi^{p}$ plus a lower order polynomial that is constrained by requiring that ${\langle{u_{p}}\,|\,{u_{p^{\prime}}}\rangle}=0$ for all $p^{\prime}<p$. The results is a sequence of orthogonal polynomials that can then be normalized according to Eq. (50).
It happens that that the weight function used to define the scalar product in Eq. (14) is of the Pollaczek-Jacobi type, but with complex valued parameters (Eqs. 12 and 17). Since our weight function may generally be evaluated using analytic continuation (see Eq. III), the Gram-Schmidt polynomials described above may be thought of as analytically continued Pollaczek-Jacobi polynomials Chen and Dai (2010). Further connections between the polynomials described above, and the Pollaczek-Jacobi polynomials will be explored in future work Foucoin:2024abc.
@@
$u_{1}(\xi)\;$ $=\;\left[{\frac{{\langle\xi^{0}\rangle}}{{\langle\xi^{0}\rangle\langle\xi^{2}\rangle-\langle\xi^{1}\rangle^{2}}}}\right]^{1/2}\;\left(\xi-\frac{\langle\xi^{1}\rangle}{\langle\xi^{0}\rangle}\right)\;.$ (51b)
In Eq. (51), we recall that $\langle\xi^{k}\rangle$ are monomial moments defined by Eq. (III). On arrives at Eqs. (51a) and (51b) regardless of whether direct or Gram-Schmidt construction is used.
@@
Since the confluent Heun polynomials are eigenfunctions to $\mathcal{D}{{}_{\xi}}$, it can be useful to represent each $|{u_{p}}\rangle$ in terms of confluent Heun polynomials, $|{y_{p^{\prime}k^{\prime}}}\rangle$. However, since the confluent Heun polynomials are overcomplete, there is no unique representation.
@@
Within Eq. (53), each canonical confluent Heun polynomial is represented in the basis of a different polynomial order, meaning that a set of multiple canonical confluent Heun polynomials would require the generation of confluent Heun polynomials from multiple orders. This is evident in Eq. (53) by $|{u_{p}}\rangle$ on the left hand side having the same label, $p$, as the upper limit of the sum on the right hand side. Thus, for each different polynomial order, a different confluent Heun polynomial basis is considered. For this reason, Eq. (53) will henceforth be referred to as the multi-order representation of the canonical confluent Heun polynomials.
@@
We emphasize that, unlike Eq. (53), $|{u_{p}}\rangle$ in the left hand side of Eq. (54) has a label $p$ that differs from the upper limit in the sum on the right hand side of Eq. (54). In Eqs. (54a) and (54b), the identity operator for the order-$p^{\prime}$ confluent Heun polynomials has been applied to $|{u_{p}}\rangle$, and in Eq. (54c),
$d_{p^{\prime}pk}\;=\;{\langle{y_{p^{\prime}k}}\,|\,{u_{p}}\rangle}\;.$ (55)
In practice, one may first generate $|{u_{p}}\rangle$ via Gram-Schmidt construction, and then explicitly compute $\langle{y_{p^{\prime}k}}\,|\,{u_{p}}\rangle$} using Eq. (19). Equation (54) will henceforth be referred to as the single-order representation of the canonical confluent Heun polynomials. If the largest polynomial order of interest is $p^{\prime}$, then the single-order representation has the practical advantage of only requiring $p^{\prime}+1$ confluent Heun polynomials to be computed after Gram-Schmidt construction.